A home bakery company sells donuts in a box that has a varied taste. From the data it turns out that the sales form a uniform discrete distribution with a maximum sales of 100 units and a minimum of 40 units per day. Forms of discrete uniform functions as follows: F(X)-|b for a SX b + 1 +1 others The question is If used random number with data a 77, m -127, Zo 12357, calculate 5 times to know the sale?
Answer #1

the random number generator given is

Zn+1=aZn mod m where a=77, m=127 and the first value for n=0 is Zo 12357

We will generate 5 random numbers as below

Zo 12357

  • Z1 = 77 × Zo mod 127-77 × 12357 mod 127 = 5
  • Zo = 77 × Z1 mod 127-77 × 5 mod 127 = 4
  • Z 77 × Z2 mod 127= 77 × 4 mod 127= 54
  • 77 × Z3 mod 127 = 77 × 54 mod 127-94
  • Z5-77x Z4 mod 127 77 × 94 mod 127-126

To convert these numbers in the interval (0,1) we divide these by m=127

  • 111-5/127 = 0.0394
  • u2-4/127 0.03 15
  • из 54/127= 0.4252
  • u_{4}=94/127=0.7402
  • u_{5}=126/127=0.9921

Next to generate number from discrete uniform distribution in the interval (a,b+1) we do the following

  • Generate a uniform random number u in the interval (0,1)
  • pick x from the interval (a,b+1) such that (x-a)/(b-a+1) le u< (x-a+1)/(b-a+1)
    • That is pick x such that a-1+(b-a+1) imes u <xle a+(b-a+1) imes u
  • here a=40 and b=100, hence we will pick an integer x such that
    • 40-1+(100-40+1) imes u <xle 40+(100-40+1) imes uimplies 39+61ule x<40+61u

the resulting xs are the sales

These are 5 draws for sales

u_1=0.0394 implies39+61 imes0.0394 le x<40+61 imes 0.0394 implies 41.402le x<42.402 implies x_1=42

U2 = 0.03 15 39 + 61 × 0.0315 < x < 40 + 61 × 0.03 15 40.921 r < 41.921 241

u_3=0.4252 implies39+61 imes0.4252 le x<40+61 imes 0.4252 implies 64.937le x<65.937 implies x_3=65

u_4=0.7402 implies39+61 imes0.7402 le x<40+61 imes 0.7402 implies 84.150le x<85.150 implies x_4=85

u_5=0.9921 implies39+61 imes0.9921 le x<40+61 imes 0.9921 implies 99.520le x<100.520 implies x_5=100

The average sale (expected sales) is

ar{x}=rac{sum x}{5}=rac{42+ 41+ 65+ 85+ 100}{5}=66.6